# Fine-Tuning Masked Diffusion for Provable Self-Correction

### v3

@@ 3.1 Theoretical foundation of PRISM
Marginalization perspective on MDM training. We rewrite the MDM training loss in equation 1.
@@
Step (a) defines the same joint distribution over $({\mathbf{x}},{\mathbf{z}})$ as MDM. Step (b) yields a random variable $({\mathbf{y}},i)$ conditioned on $\mathbf{z}$}. Although we do not designate a sampling rule for the index $i$, PRISM’s guarantee is invariant to its rule, e.g., uniformly random, confidence-based. As a result, PRISM induces a joint distribution over $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$. Taking expectation over this joint distribution, we define the PRISM loss:
@@
Proposition 1 (PRISM).
The per-token quality uniquely minimizes the PRISM loss $\mathcal{L}(\phi)$ equation 3: $g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}).$
Before proving Proposition 1, we highlight two key advantages that PRISM offers.
@@
PRISM: Guarantee. Proposition 1 can be demonstrated within a few lines. The key mechanism is that the joint distribution $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ exhibits a simple posterior given $({\mathbf{y}},i)$; $\mathbf{z}$} is deterministically ${\mathbf{z}}={\mathbf{y}}\oplus\mathbf{m}_{i}$ and each ${\mathbf{x}}^{i}$ is drawn from the same unmasking posterior of MDM, i.e., $v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$. Therefore, fixing $({\mathbf{y}},i)$ and then marginalizing the binary label $\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]$ over $({\mathbf{x}},{\mathbf{z}})$ yields the simple byproduct,
@@
which is exactly minimized at $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}.
@@ 3.2 Empirical design choice of PRISM
Figure 2: PRISM training pipeline. Fine-tuning samples are built in two steps; (a) masking process to obtain $({\mathbf{x}},{\mathbf{z}})$ and (b) unmask a few indices in $\mathbf{z}$} to get $\mathbf{y}$} with a pretrained MDM $f_{\theta}$ (Left). We add a lightweight adapter to a pretrained MDM so that the resulting model computes the unmasking posterior ($f_{\theta}$) and per-token quality ($g_{\theta}$) (Right).
@@
Plug-and-Play approach. Assume a pretrained MDM $f_{\theta}$ is given, which by design has learned the unmasking posterior. PRISM augments this model with a lightweight auxiliary adapter that reuses the same backbone to produce a size-$L$ tensor. In words, the unmasking posterior and per-token quality share a single backbone and are computed by separate heads (Figure 2).
@@
For a given sequence, we respectively denote the two head outputs as $f_{\theta}$ and $g_{\theta}$. Let $\mathbf{h}_{\theta}$ denote the final hidden state of the MDM backbone. The unmasking posterior $f_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the unmasking head (which originally exists in pretrained MDM) and then applying the softmax. The per-token quality $g_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the attached head and applying the sigmoid. Sigmoid is used to guarantee $g_{\theta}^{i}\in[0,1]$, making it suitable for the binary cross-entropy loss. Thus, with a single $\theta$ and a lightweight adapter, we jointly model the unmasking posterior and then the per-token quality: for a partially masked sequence $\mathbf{w}$},
@@
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{w}})}\approx p({\mathbf{x}}^{i}={\mathbf{w}}^{i}\,|\,{\mathbf{w}}\oplus\mathbf{m}_{i})\in[0,1],\quad{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathbf{w}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{w}})\in\Delta(\mathcal{V}).$
At a high level, the PRISM recipe is simple: train $g_{\theta}$ using the PRISM loss defined in equation 3. Consequently, the parameters updated through $g_{\theta}$ are those of the attached head and the backbone, excluding the unmasking head. As an alternative to fine-tuning the backbone, one can add a LoRA adapter (hu2022lora) into the backbone and train only the LoRA adapter parameters.
Practical interventions. We introduce two practical interventions that improve PRISM’s efficiency. First, to preserve $f_{\theta}$’s ability to estimate the unmasking posterior while $g_{\theta}$ is learning toward the per-token quality, we add the MDM training loss equation 3 as a regularization term. For a regularization constant $\lambda>0$, the total loss is therefore:
$\mathcal{L}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i)}\Bigg[{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}\,)}+\lambda\times\frac{1}{n}\sum_{j\colon{\mathbf{z}}^{j}=\mathbf{m}}{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}-\log f_{\theta}^{j}({\mathbf{x}}^{j}\,|\,{\mathbf{z}})}\Bigg].$ (4)
Second, in the sampling process for constructing PRISM training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$, we modify step (b). To obtain $\mathbf{y}$}, rather than unmasking a single $\mathbf{y}$} from $\mathbf{z}$}, we obtain multiple $({\mathbf{y}},i)$ pairs from one $\mathbf{z}$}. This produces multiple training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ with varying $({\mathbf{y}},i)$ values from a single forward pass of $f_{\theta}(\cdot\,|\,{\mathbf{z}})$, improving data efficiency. Since this procedure of sample construction minimizes $g_{\theta}$ (not $f_{\theta}$), we apply stop-gradient to this $f_{\theta}$. The pseudocode is given as Algorithm 1.
Algorithm 1 PRISM fine-tuning from a pretrained MDM
1:Require: Pretrained MDM backbone $f_{\theta}$, per-token quality head $g_{\theta}$, optimizer $\mathcal{O}$, true distribution $p_{\mathrm{data}}$, regularization constant $\lambda>0$, number of unmasking tokens $k\in\mathbb{N}$.
3:Initialize the loss $\mathcal{L}(\theta)\leftarrow 0$
[…]
5:  # Sample $\mathbf{y}$} from $\mathbf{z}$} with $f_{\mathrm{sg}(\theta)}$
6:  Select a subset $\mathcal{S}\subseteq\{i\mid{\mathbf{z}}^{i}=\mathbf{m}\}$ with $|\mathcal{S}|=k$.
@@
7:  Obtain $\mathbf{y}$} by replacing ${\mathbf{z}}^{i}=\mathbf{m}$ to $v^{i}\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}f_{\mathrm{sg}(\theta)}^{i}(\cdot\,|\,{\mathbf{z}})}$ for each $i\in\mathcal{S}$.
8:  # Calculate Loss
9:  $\mathcal{L}(\theta)\leftarrow\mathcal{L}(\theta)+\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}{\mathrm{BCE}(\,\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]\,,\,{{\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathbf{y}})}}\,)}$ $\triangleright$ PRISM Loss
10:  $\mathcal{L}(\theta)\leftarrow\mathcal{L}(\theta)+\lambda\times\frac{1}{n}\sum_{j:{\mathbf{z}}^{j}=\mathbf{m}}{\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}-\log f_{\theta}^{j}({\mathbf{x}}^{j}\,|\,{\mathbf{z}})}$. $\triangleright$ Regularization (MDM loss)
11:  $\theta\leftarrow\mathcal{O}(\theta,\mathcal{L}(\theta))$
12:end while
@@
Importantly, the theoretical guarantee in Proposition 1 still maintains: the true minimizer $g_{\theta}$ of this modified PRISM loss continues to encode meaningful per-token quality. We defer the precise statement and proof to Appendix B.3. Moreover, we conduct a careful ablation of different design choices on sampling $\mathbf{y}$} in Appendix D.3 and Appendix D.4.
@@ 4 Experiments
In this section, we demonstrate that PRISM is a practically efficient and scalable approach across different domains and scales. In Section 4.1, we apply PRISM to a 3030M-scale MDM for Sudoku. In Section 4.2, we apply PRISM to a 170M-scale MDM and evaluate its ability for natural language modeling. Finally, in Section 4.3, we apply PRISM to LLaDA-8B-Instruct (nie2025large) and evaluate it on a coding task. We first outline the pipeline shared across all experiments. We provide the experimental details in Appendix D.
Experiment pipeline. We start from an MDM pretrained for each domain and attach a lightweight adapter. We then fine-tune the model on the PRISM loss defined in equation 4. For baselines, we adopt ReMDM (ReMDM-cap or ReMDM-loop) and ReMDM-conf from wang2025remasking; for fair comparison, we re-evaluate these baselines using the same backbone after PRISM fine-tuning. Finally, since our focus is on the remasking strategy, we restrict comparisons to remasking baselines. We summarize our experimental results as follows:
@@ 4.1 Motivating example: Sudoku puzzle
Setup. We construct a corpus of Sudoku puzzles (4848k for training and 22k for evaluation), following the setup of (ben2025accelerated; alp2024sudoku). Each board is tokenized as a length-$89$ sequence comprising $81$ cell values and $8$ [EOL] tokens that separate columns. Following the MDM pretraining recipe of sahoo2024simple, we use a 28.628.6M Diffusion Transformer (DiT; peebles2023scalable) architecture as the backbone and pretrain it for 100100k iterations ($\approx 530$ epochs). We then fine-tune it with the PRISM loss, additionally with a lightweight adapter (0.130.13M parameters).
Figure 3: Left: PRISM achieves higher success accuracy on Sudoku puzzles than baselines. (red line) Right: PRISM detects the incorrect cells by assigning low per-token quality (red-colored cells).
Results. We evaluate the success rate for the same model across epochs for PRISM and baselines. As shown in Figure 3, PRISM achieves progressively higher success rates during fine-tuning (red line) and surpasses the baselines. This confirms that the learned per-token quality enables effective self-correction at inference. Notably, PRISM starts to outperform vanilla MDM inference (dashed line) and baselines within only a few epochs, far fewer than the pretraining epochs ($\approx 530$), corroborating our claim of sample efficiency.
@@ 4.2 Unconditional Text Generation
Setup. We first take a DiT-based MDM with 170170M parameters, using a checkpoint from (sahoo2024simple). This MDM is pretrained on the OpenWebText corpus (Gokaslan2019OpenWeb) with the GPT-2 tokenizer (radford2019language), a maximum sequence length $L=1024$, and a total of 262262B tokens. We then attach a lightweight 77M parameter adapter, one attention layer applied to the final hidden state, and fine-tune it with the PRISM loss on a total 164164M tokens from the same OpenWebText corpus. This fine-tuning is data-efficient, using $\mathbf{1600}\times$ fewer tokens than pretraining.
@@
Results. We then assess the performance of PRISM and baselines on unconditional generation. Our primary metrics are Generative Perplexity and MAUVE (pillutla2021mauve), which respectively assess per-sequence fluency and distributional similarity between generated samples and the reference distribution. Moreover, given the prior observation (zheng2024masked) that a sequence with too many redundant tokens can spuriously yield better (i.e., lower) generative perplexity, we also report the entropy as a sanity metric. Following standard evaluation protocols (pillutla2021mauve; wang2025remasking), we generate $5000$ samples to compute both metrics, using GPT-2-Large as the reference model. As shown in Figure 4, PRISM outperforms baselines on both metrics (red lines) while achieving a comparable entropy, especially in the regime with fewer sampling steps ($N<1024$).
Sampling strategy. As sentence diversity is crucial for unconditional generation, we observe that the vanilla PRISM inference from Section 3.3 can degrade diversity. We attribute this to the greedy nature of the token-remasking policy that ranks positions by per-token quality. To mitigate this, we adopt PRISM-loop, analogous to ReMDM-loop (wang2025remasking): follow PRISM’s remasking scheme and add loop steps at $t_{N}=0$. After obtaining a clean sequence, in a loop, we randomly remask a subset of tokens and unmask them. A full breakdown of the sampling strategy is given in Appendix C.1,  D.2.
Figure 4: Unconditional text generation performance. Metrics are evaluated at $\text{NFE}\in\{128,256,512,1024,2048,4096\}$; black dashed lines denote validation set references. PRISM (red) outperforms baselines (ReMDM: green, ReMDM-conf: blue, Vanilla MDM: gray), particularly at lower sampling steps ($N<1024$). Detailed numerical results are reported in Table 3.
Quantitative analysis. To further investigate PRISM’s behavior at a mechanistic level, we evaluate how well its per-token quality scores are calibrated to the desired likelihood. For partially masked sequences $\mathbf{x}_{t}$ encountered at inference, we collect per-token quality scores $g_{\phi}^{i}(\mathbf{x}_{t})$ over all clean positions $i$ and group these tokens into buckets according to the predicted scores. For each bucket, we then estimate an empirical oracle likelihood and query it by a pretrained model $\theta$; $f_{\theta}^{i}(\mathbf{x}_{t}^{i}\mid\mathbf{x}_{t}\oplus\mathbf{m}_{i})$ for each $i$. Interestingly, the resulting calibration curves (Figure 7) closely follow the queried likelihoods, indicating that $g_{\theta}^{i}(\mathbf{x}_{t})$ is a well-calibrated estimate of the target posterior, as desired.
4.3 Appyling PRISM to 88B MDM

### v4

@@ 3.1 Theoretical Foundation of PRISM
Marginalization perspective on MDM training. We rewrite the MDM training loss in (1).
@@
Step (a) defines the same joint distribution over $({\mathbf{x}},{\mathbf{z}})$ as MDM. Step (b) yields a random variable $({\mathbf{y}},i)$ conditioned on $\mathbf{z}$}. Although we do not designate a sampling rule for the index $i$, PRISM’s guarantee is invariant to its rule, e.g., uniformly random, confidence-based. As a result, PRISM induces a joint distribution over $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$; for illustration, please see Figure 1 (right). Taking expectation over this joint distribution, we define the PRISM loss:
@@
Proposition 3.1 (PRISM).
The per-token quality uniquely minimizes the PRISM loss $\mathcal{L}(\phi)$ (2): $g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i}).$
Before proving Proposition 3.1, we highlight two key advantages that PRISM offers.
@@
PRISM: Guarantee. Proposition 3.1 can be proven within a few lines. The key mechanism is that the joint distribution $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ exhibits a simple posterior given $({\mathbf{y}},i)$; $\mathbf{z}$} is deterministically ${\mathbf{z}}={\mathbf{y}}\oplus\mathbf{m}_{i}$ and each ${\mathbf{x}}^{i}$ is drawn from the same unmasking posterior of MDM, i.e., $v\sim p({\mathbf{x}}^{i}=\cdot\,|\,{\mathbf{z}})$. Therefore, fixing $({\mathbf{y}},i)$ and then marginalizing the binary label $\mathbf{1}[{\mathbf{x}}^{i}={\mathbf{y}}^{i}]$ over $({\mathbf{x}},{\mathbf{z}})$ yields the simple byproduct,
@@
which is minimized at $\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\phi^{\star}}^{i}({\mathbf{y}})=p({\mathbf{x}}^{i}={\mathbf{y}}^{i}\,|\,{\mathbf{y}}\oplus\mathbf{m}_{i})$}.
@@ 3.2 Empirical Design Choice of PRISM
Plug-and-Play approach. Assume a pretrained MDM $f_{\theta}$ is given, which by design has learned the unmasking posterior. PRISM augments this model with a lightweight auxiliary adapter that reuses the same backbone to produce a size-$L$ tensor. In words, the unmasking posterior and per-token quality share a single backbone and are computed by separate heads.
@@
For a given sequence, we respectively denote the two head outputs as $f_{\theta}$ and $g_{\theta}$. Let $\mathbf{h}_{\theta}$ denote the final hidden state of the MDM backbone. The unmasking posterior $f_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the unmasking head (which originally exists in pretrained MDM) and then applying the softmax. The per-token quality $g_{\theta}$ is obtained by passing $\mathbf{h}_{\theta}$ through the attached head and applying the sigmoid. Sigmoid is used to guarantee $g_{\theta}^{i}\in[0,1]$, making it suitable for the binary cross-entropy loss. Thus, with a single $\theta$ and a lightweight adapter, we jointly model the unmasking posterior and then the per-token quality: for a partially masked sequence $\mathrm{w}$},
@@
${\color[rgb]{0,0.625,0.52734375}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.625,0.52734375}g_{\theta}^{i}({\mathrm{w}})}\approx p({\mathbf{x}}^{i}={\mathrm{w}}^{i}\,|\,{\mathrm{w}}\oplus\mathbf{m}_{i})\in[0,1],$
@@
${\color[rgb]{0.25390625,0.41015625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.25390625,0.41015625,0.8828125}f_{\theta}^{i}(\cdot\,|\,{\mathrm{w}})}\approx p({\mathbf{x}}^{i}=\cdot\,|\,{\mathrm{w}})\in\Delta(\mathcal{V}).$
At a high level, the PRISM recipe is simple: train $g_{\theta}$ using the PRISM loss defined in (2). Consequently, the parameters updated through $g_{\theta}$ are those of the attached head and the backbone, excluding the unmasking head. As an alternative to fine-tuning the backbone, one can add a LoRA adapter (Hu et al., 2022) into the backbone and train only the adapter parameters.
Practical interventions. We introduce two practical interventions that improve PRISM’s efficiency. First, to preserve $f_{\theta}$’s ability to estimate the unmasking posterior while $g_{\theta}$ is learning toward the per-token quality, we add the MDM training loss (1) as a regularization term.
Second, in the sampling process for constructing PRISM training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$, we modify step (b). To obtain $\mathbf{y}$}, rather than unmasking a single $\mathbf{y}$} from $\mathbf{z}$}, we obtain multiple $({\mathbf{y}},i)$ pairs from one $\mathbf{z}$}. This produces multiple training pairs $({\mathbf{x}},{\mathbf{z}},({\mathbf{y}},i))$ with varying $({\mathbf{y}},i)$ values from a single forward pass of $f_{\theta}(\cdot\,|\,{\mathbf{z}})$, improving data efficiency. Since this procedure of sample construction minimizes $g_{\theta}$ (not $f_{\theta}$), we apply stop-gradient to this $f_{\theta}$. We also provide the pseudocode at Algorithm 1.
@@
Importantly, the theoretical guarantee in Proposition 3.1 still maintains: the true minimizer $g_{\theta}$ of this modified PRISM loss continues to encode meaningful per-token quality. We defer the precise statement and proof to Appendix B.3. Moreover, we conduct a careful ablation of different design choices on sampling $\mathbf{y}$} in Appendix D.3 and Appendix D.4.
@@ 4 Experiments
In this section, we demonstrate that PRISM is a practically efficient and scalable approach across different domains and scales. In Section 4.1, we apply PRISM to a 3030M-scale MDM for Sudoku. In Section 4.2, we apply PRISM to a 170M-scale MDM and evaluate its ability for natural language modeling, and in Section 4.3, we apply PRISM to LLaDA-8B-Instruct (Nie et al., 2025) and evaluate it on a Python coding task. In Section 4.4, we conduct a mechanistic analysis on PRISM to further showcase its efficiency. We first outline the pipeline shared across all experiments. Experiment pipeline. Starting from a pretrained MDM, we attach a lightweight adapter and fine-tune it with the PRISM loss. We compare against remasking baselines ReMDM (ReMDM-cap/ReMDM-loop) and ReMDM-conf (Wang et al., 2025), re-evaluated using the same backbone after PRISM fine-tuning. We summarize our experimental results as follows:
@@
  • Operation: PRISM fine-tuned MDMs calibrate the per-token qualities as desired, and qualitatively correct errors in coding problems.
@@ 4.1 Motivating Example: Sudoku Puzzle
Figure 2: Left: PRISM achieves higher accuracy on Sudoku puzzles than baselines. (red curve) Right: PRISM detects the incorrect cells by assigning low per-token quality (red-colored cells).
Setup. We construct a corpus of Sudoku puzzles (4848k for training and 22k for evaluation), following the setup of (Ben-Hamu et al., 2025; Alp, 2024). Each board is tokenized as a length-$89$ sequence comprising $81$ cell values and $8$ [EOL] tokens that separate columns. Following the MDM pretraining recipe of Sahoo et al. (2024), we use a 28.628.6M Diffusion Transformer (DiT; Peebles and Xie (2023)) architecture as the backbone and pretrain it for 100100k iterations ($\approx 530$ epochs). We then fine-tune it with the PRISM loss, additionally with a lightweight adapter (0.130.13M parameters).
Results. We evaluate the success rate for the same model across epochs for PRISM and baselines. As shown in Figure 2, PRISM achieves progressively higher success rates during fine-tuning (red curve) and surpasses the baselines. This confirms that the learned per-token quality enables effective self-correction at inference. Notably, PRISM starts to outperform vanilla MDM inference (dashed line) and baselines within only a few epochs, far fewer than the pretraining epochs ($\approx 530$), corroborating our claim of sample efficiency.
@@
Figure 3: (a) Unconditional text generation performance. Metrics are evaluated at $\text{NFE}\in\{64,128,256,512,1024\}$; PRISM (red) outperforms baselines (ReMDM: green, ReMDM-conf: blue, Vanilla MDM (without remasking): gray), particularly at lower sampling steps. Detailed numerical results are reported in Table 2. (b) Calibration study. PRISM fine-tuned model’s per-token quality closely tracks the empirical likelihood.
@@ 4.2 Unconditional Text Generation
Setup. We first take a DiT-based MDM with 170170M parameters, using a checkpoint from (Sahoo et al., 2024). This MDM is pretrained on the OpenWebText corpus (Gokaslan and Cohen, 2019) with the GPT-2 tokenizer (Radford et al., 2019), a maximum sequence length $L=1024$, and a total of 262262B tokens. We then attach a lightweight 77M parameter adapter, one attention layer applied to the final hidden state, and fine-tune it with the PRISM loss on a total 164164M tokens from the same OpenWebText corpus. This fine-tuning is data-efficient, using $\mathbf{1600}\times$ fewer tokens than pretraining. During fine-tuning, we select the subset $\mathcal{S}$ based on confidence scores.
@@
Results. We then assess the performance of PRISM and baselines on unconditional generation. Our primary metrics are Generative Perplexity and MAUVE (Pillutla et al., 2021), which respectively assess per-sequence fluency and distributional similarity between generated samples and the reference distribution. Moreover, given the prior observation (Zheng et al., 2024) that a sequence with too many redundant tokens can spuriously yield better, i.e., lower generative perplexity, we report the entropy as a sanity metric. Following standard evaluation protocols (Pillutla et al., 2021; Wang et al., 2025), we generate $5000$ samples to compute both metrics, using GPT-2-Large as the reference model. As shown in Figure 3.(a), PRISM outperforms baselines on both metrics (red curves) while achieving a comparable entropy, especially in the regime with fewer sampling steps. See Appendix D.1 for more details.
HumanEval MBPP
Sampling steps 256 512 1024 256 512 1024
PRISM 28.0 39.0 42.7 21.8 29.1 32.3
ReMDM 26.2 34.8 42.5 19.7 28.5 32.9
ReMDM-conf 25.6 34.1 36.6 20.1 29.0 32.5
MDM 25.6 34.1 36.6 18.2 27.8 31.9
Table 1: Zero-shot performance on coding tasks.
4.3 Applying PRISM to 8B MDM